Through careful record keeping, you have the times to burn out of 100 Bright brand light bulbs. The average time to burn out is 200 hours, with a standard deviation of 12 hours. Give a 95% confidence interval for the mean time to burn out of this brand of light bulb.


1
Expert's answer
2022-05-26T07:15:53-0400

The critical value for "\alpha = 0.05" and "df = n-1 = 99" degrees of freedom is (using critical values table)

"t_c = z_{1-\alpha/2; n-1} = 1.984"

The corresponding confidence interval is computed as shown below:



"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}},\bar{x}+t_c\times\dfrac{s}{\sqrt{n}})""=(200-1.984\times\dfrac{12}{\sqrt{100}},200+1.984\times\dfrac{12}{\sqrt{100}})""=(197.62,202.38)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "197.62 < \mu < 202.38," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "197.62, 202.38)."